# COMET: Co-Optimization of a CNN Model using Efficient-Hardware OBC Techniques

### v2

@@ Abstract
Convolutional Neural Networks (CNNs) are highly effective for computer vision and pattern recognition tasks; however, their computational intensity and reliance on hardware such as FPGAs pose challenges for deployment on low-power edge devices. In this work, we present COMET, a framework of CNN designs that employ efficient hardware offset-binary coding (OBC) techniques to enable co-optimization of performance and resource utilization. The approach formulates CNN inference with OBC representations of inputs (Scheme A) and weights (Scheme B) separately, enabling exploitation of bit-width asymmetry. The shift–accumulate operation is modified by incorporating the offset term with the pre-scaled bias. Leveraging inherent symmetries in Schemes A and B, we introduce four novel look-up table (LUT) techniques—parallel, shared, split, and hybrid—and analyze them to identify the most efficient options. Building on this foundation, we develop an OBC-based general matrix multiplication core using the im2col transformation, enabling efficient acceleration of a fixed-point modified LeNet-5 model. FPGA evaluations demonstrate that the proposed co-optimization approach significantly reduces resource utilization compared to state-of-the-art LeNet-5 based CNN designs, with minimal impact on accuracy.
@@
Convolutional Neural Networks (CNN), Field-Programmable Gate Array (FPGA), General Matrix-Multiply (GEMM), Look-up Table (LUT), Offset-Binary Coding (OBC).
@@ I Introduction
Convolutional neural networks (CNNs) have become a cornerstone of modern deep learning. Their ability to learn features directly from raw image data in an end-to-end manner has enabled them to significantly surpass traditional handcrafted methods in key computer vision and pattern recognition tasks, including image classification, object detection, and semantic segmentation [1]. Owing to these advantages, CNNs have found widespread use in diverse domains such as industrial defect inspection, medical imaging, and others [2]. However, despite their success, conventional CNN inference relies on multiplication-intensive operations that are power-hungry and chip-area demanding, posing challenges for edge platforms. To address these challenges, CNNs have been increasingly implemented on FPGAs [3, 4, 5], which are particularly well-suited for CNN acceleration due to their inherent parallelism, reconfigurability, and efficient utilization of on-chip memory and digital signal processing (DSP) slices [6].
To optimize resource utilization on FPGA, many CNN accelerators adopt the im2col transformation [7] to restructure 2D convolutions into tile-based general matrix multiplication (GEMM) workloads, enabling the reuse of a shared multiply–accumulate (MAC) array across layers [8]. While this strategy improves computational throughput, it exacerbates the demand for DSP multipliers and increases pressure on on-chip memory bandwidth. To reduce both the number of DSP multipliers and memory bandwidth, various reduction techniques have been explored, including Winograd minimal filtering [9], sparse convolution [10], and depthwise separable convolutions [11]. Although these methods effectively reduce the number of operations, they remain tied to DSP-centric architectures—limitations that are particularly pronounced in edge scenarios where both DSP availability and dynamic power budgets are highly constrained.
For efficient CNN implementation on resource-constrained hardware, asymmetric quantization is particularly beneficial, as reducing the weight bit-width (while keeping the input bit-width fixed) can significantly lower power and energy consumption, albeit at the expense of some accuracy [12]. To effectively evaluate such methods, it is essential to select representative models that balance computational complexity and accuracy. Lightweight architectures are particularly well-suited for exploring low-area, low-power design strategies, especially where minimizing energy consumption and silicon footprint is critical. While several works  [13, 14, 4] have implemented complex models such as MobileNetV2 and SqueezeNet—leveraging depthwise separable convolutions and compact network designs to optimize throughput—these approaches often depend on MAC operations implemented through DSP blocks. However, in severely resource-constrained environments, they become inefficient due to their high computational and energy cost.
LeNet-5 model [15, 16] is often selected as a representative small-scale CNN because of its: (1) simple architecture and low parameter count, which make it suitable for resource-constrained hardware [17]; (2) well-established MNIST performance, providing a reproducible benchmarking baseline [18]; and (3) compatibility with operator types and problem scales typically targeted by GEMM-based frameworks [19]. Unlike the aforementioned models,Its simple architecture enables a precise study of the implementation trade-offs. Following this rationale, we adopt the classic LeNet-5 CNN as our evaluation benchmark. The model is compact enough to fit entirely on mid-range FPGAs, yet still encompasses representative convolutional, pooling, and fully connected layers. Recent studies [20, 21] indicate that while architectural choices strongly influence model accuracy, their effects are often intertwined with training factors. For example, replacing tanh activations with ReLU has been shown to improve performance in practice [20], and substituting large fully connected layers with global average pooling (GAP) can yield a more efficient design [21].
Distributed arithmetic (DA) is a multiplier-less approach ideal for low-power, area-constrained hardware [22, 23]. It replaces MAC operations in inner-product computation (IPC) with precomputed look-up tables (LUTs) and a shift-accumulate (SA) unit. This involves representing one vector in two’s complement (TC) [24] or offset-binary coding (OBC) [25], and storing binary or offset binary combinations of the other vector in LUTs. While TC causes exponential LUT growth, OBC reduces this by half through symmetry. Recent works have optimized DA-based LSTMs using hardware-based LUTs [26, 27, 28, 29]. For instance, [26] proposed a parallel OBC-based LUT with pipelining, which reduced critical path delay (CPD) but increased hardware complexity. While [28] partially overcomes the limitations of [26] by adopting a TC-based approach, some challenges in optimizing hardware cost remain unaddressed. Furthermore, serial LUT designs using high-radix OBC and TC were proposed in [27, 29], incurring additional clock cycle latency. The serial-LUT suffers from limited reusability and significant delay for large dimensions, making pipelining essential but adding latency. To the best of our knowledge, only a few DA-specific implementations targeting either 1-point convolution or CNN acceleration (on ASICs) exist [30, 31, 32]; for example, [30] minimized the LUT size via input pairing, while [31] reduced memory usage through in-memory optimization.
@@
Although various algorithmic optimizations have helped reduce MAC operation counts in CNNs, they still fundamentally rely on DSP-based multipliers, which remain a critical bottleneck for low-power edge deployments. Even advanced dataflow strategies [11, 3] that aim to maximize data reuse fail to eliminate this reliance on MAC units, thereby limiting the efficiency and scalability of accelerators targeting stringent energy and area constraints. This persistent challenge motivates the design of a fully DSP-free CNN architecture based on LeNet-5 model. In this work, we first adopt a streaming (im2col + tiled) strategy to partition the original feature map into multiple small matrix blocks and perform GEMM on each block, thereby avoiding IPCs on large vectors in a single pass. To eliminate large pre-computed LUTs, we propose four dynamic parallel LUT structures that enable flexible sharing and reuse via runtime computation of the partial sums. A CNN architecture relying on the OBC scheme is developed and integrated into a general-purpose GEMM core. Using im2col, standard CNN kernels are mapped and executed as OBC-accelerated GEMM operations. We name resulting framework COMET (Co-Optimization of a CNN Model using Efficient Hardware OBC Techniques), a framework for efficient CNN acceleration that leverages hardware-optimized OBC techniques to co-optimize performance and resource utilization. Our main contributions are as follows:
  • We propose a compact im2col-based formulation of the 2D CNN convolution layer using OBC representations of inputs (Scheme A) and weights (Scheme B), enabling systematic analysis for hardware optimization.
  • We introduce, as part of the COMET framework, four novel LUT architectures based on efficient hardware OBC techniques. We enhance the shift–accumulate unit to merge pre-scaled bias and offset terms, thereby enabling co-optimization of resource and energy efficiency across multiple clock frequencies.
  • We map LeNet-5 CNN kernels onto an OBC-GEMM framework with unified accumulation using Scheme A and Scheme B, and we analyze the impact of quantization on model accuracy.
  • We explore, implement, and evaluate the design space of the proposed architectures on an FPGA platform, providing a comprehensive analysis of design trade-offs such as resource utilization and energy efficiency.
The remainder of this paper is organized as follows. Section II introduces the OBC formulation of a standard CNN. Section III presents the proposed LUT architectures and analyzes their co-optimization trade-offs. Section IV discusses the design considerations and evaluation of the proposed OBC-GEMM module. Section V provides a comparative analysis of the proposed design against state-of-the-art CNN accelerators. Finally, Section VI concludes the paper.
Figure 2: OBC-based IPC for $K=4$ with (a) traditional LUT and SA-unit [25], (b) hardware LUT (w/o pipelining) and CSFA-based SA unit [26].
@@ II CNN Formulation with OBC
As shown in Fig. 1, a standard CNN is a multi-layer pipeline in which each layer transforms input feature maps of size $H\times W$ into higher-level representations using convolutional kernels of size $K\times L$ and biases. A generic layer has $C$ input channels and $N$ output channels. Each output channel is obtained by convolving all $C$ input channels with their corresponding kernels via a sliding-window operation (A), summing across channels, and repeating this process for all $N$ outputs. For clarity, we consider the case of batch size $1$, although the formulation extends naturally to multiple batches. Specifically, given an input tensor with $C$ channels, each with spatial dimensions $H\times W$, i.e., $X\in\mathbb{R}^{C\times H\times W}$, the output of a 2D convolution of $X$ with a filter tensor $\Theta\in\mathbb{R}^{N\times C\times K\times L}$ is a tensor $Y\in\mathbb{R}^{N\times(H+K)\times(W+L)}$, whose entries are given by
@@
where $n=1,\ldots,N$, $h=1,\ldots,H+K$, $w=1,\ldots,W+L$, and $\beta_{n}\in\mathbb{R}$ represents a bias term. Note  (1) describes the convolution computation for a single layer. Each product $X_{c,h+k,w+l}\Theta_{n,c,k,l}$ followed by its accumulation constitutes one MAC operation. Therefore, a single convolution layer requires approximately $N\times H\times W\times C\times K\times L$ MAC operations. Across multiple layers, the total MACs can reach tens of millions, imposing heavy demands on compute and memory bandwidth [13]. To address this issue, the im2col transformation [7] enables tile-based GEMM, allowing a single MAC array to be reused across layers. Applied to (1), the transformation generally follows these steps:
@@
  • The convolution is expressed as an inner product of length $N_{p}=CKL$ between the flattened input patch and filter:
@@
The resulting output is then reshaped into the appropriate order. For consistency in IPC formulation, we set $C=1$ and $L=1$ for simplicity, while the analysis can be extended for any $C$ and $L$. By denoting the summation in (2) as $\tilde{Y}$}, we have
@@
where $x_{i,r}\in\{0,1\}$ is the $r$th bit of $x_{i}$. In OBC representation, $x_{i}$ can be expressed as $x_{i}=\frac{1}{2}[x_{i}-(-x_{i})]$, which results (4) in
@@
where $\bar{x}_{i,r}$ is the ones complement of $x_{i,r}$ and $\lfloor\cdot\rfloor$ is the floor function. Substituting (5) and (6) into (3) and rearranging, we have
@@
Each LUT entry corresponds to a possible combination of weights with input bit-slices $(x_{n,B_{1}-1},\ldots,x_{n,0};\ 1\leq n\leq K)$, ordered from least to most significant bit. During runtime, these bit-slices serve as addresses to access the LUT. For each address access, the output is accumulated using a SA operation, repeated over $B_{1}$ clock cycles to compute the output $\tilde{Y}$, as shown in Fig. 2(a). Upon closer observation, the offset term $\theta_{ot}$ corresponds to 2’s complement of the LUT content stored at the $0^{\textnormal{th}}$ address location. This value is accessed when loading the offset term at the beginning of shift-accumulation, as illustrated in Fig. 2(a). While traditional LUTs can store weight combinations, they become inefficient for large $K$ due to increased access time. An alternative in [26] uses adders and multiplexers to generate combinations dynamically but lacks scalability and introduces significant hardware overhead. Specifically, implementing such a LUT requires $2^{K}(1-2^{-K/2})$ adders/subtractors and $(2^{K-1}-1)$ 2-to-1 multiplexers. The CPD is $\log_{2}K\cdot T_{A}+(K-1)\cdot T_{MX}$, where $T_{A}$ and $T_{MX}$ are the delays of an adder and a multiplexer, respectively. To mitigate this, carry-save full adders (CSFAs) have been used to reduce delay.
Similar to (11), an expression for $Y$ can be obtained with Scheme B, where each weight $\theta_{i}$, represented in OBC form with a bit-width of $B_{2}$, follows the steps indicated by (4)–(11). In this case, we would have
@@
Note that Scheme B has the same structure as Scheme A (Fig. 2), except that the LUT content and the number of clock cycles in the SA unit may differ ($B_{1}\neq B_{2}$). These schemes differ only in choosing which component to represent in OBC form, exploiting the asymmetry between weight and input bit-widths to provide hardware implementation trade-offs.
@@ III-A System Overview
The proposed CNN accelerator is implemented on the Xilinx ZCU106 platform. All input feature maps, weights, and biases are pre-loaded into on-chip RAM before execution, and the computed outputs are written back to the same memory. To favor edge deployment and design flexibility, all buffers are implemented using Slice RAM rather than dedicated BRAM. The accelerator design workflow consists of two stages: (1) model formulation with hardware-oriented optimization, and (2) deployment on the FPGA. The CNN model is first expressed using the two OBC schemes (A and B), which define the GEMM decomposition strategy, and is then mapped to hardware-optimized architectures.
@@ III-B Address Generator and Control Unit
At the core of the address generator is a set of nested counters operating under the FSM’s supervision, responsible for address generation and read/write data scheduling. As shown in Fig. 4, the counter architecture consists of a dedicated cntr0 and three hierarchical counter groups for read (rd_cntr#$m$), calculation (cal_cntr#$m$), and write (wr_cntr#$m$) operations. Each group contains four levels of counters indexed by $m\in\{1,4\}$, corresponding respectively to tiling (i.e., partitioning a convolution into smaller blocks for efficient computation), spatial position $(h,w)$ as defined in (1), output channel index $(n)$, and layer index. The standalone cntr0 tracks the clock cycle index within the OBC-GEMM computation. The signals marked 1 to 4 in Fig. 4 represent the hierarchical carry mechanism, which enables stepwise progression within each counter group and governs the transfer of counter values between the read, calculation, and write stages.
In this hierarchical carry mechanism, hierarchical counter rippling applies primarily to the read counter stage. The clock cycle index counter (cntr0) generates sequential memory addresses for data fetches within one tile-level OBC-GEMM operation. When it reaches its upper bound (equal to the number of fetch cycles per tile), all data for the current tile has been loaded. At this point, carry signal 1 is asserted, cntr0 resets, and rd_cntr1 (tile index) increments to begin loading the next tile. When rd_cntr1 completes its range (number of tiles per spatial location), carry 2 is raised, rd_cntr1 resets, and rd_cntr2 (spatial position) increments. Similarly, once rd_cntr2 reaches its bound (number of spatial positions per output channel), carry 3 triggers, resetting rd_cntr2 and incrementing rd_cntr3 (output channel index). Finally, when rd_cntr3 reaches the total output channels for the layer, carry 4 asserts, resetting rd_cntr3 and incrementing rd_cntr4 (layer index) to start the next layer.
@@
[…]
III-C OBC-GEMM Core
The OBC-GEMM core consists of an array of $L$ LUTs each of size $K$, $L$ SA units, and $L$ PISO units each of size $K$ that handle bit-slices of inputs (or weights), as shown in Fig. 5. The LUTs dynamically generate partial sums using adders and multiplexers for hardware acceleration of IPCs, while SA units efficiently combine these results. PISO units bit-serialize the inputs (or weights) to fit the data flow of the scheme A (or scheme B). A key optimization is a single dedicated adder that computes the offset and bias terms. This eliminates separate bias computation, reducing hardware complexity, and saving area and power.
@@ III-C1 Parallel LUT
III-C2 Shared LUT
As the name suggests, the shared LUT approach is derived from the parallel LUT by sharing the generated LUT contents to reduce the generated partial sums, as illustrated in Fig. 6(a). Specifically, it is observed that the LUT contents at address locations 4–7 can be derived by simply loading $\theta_{2}$ instead of $\theta_{2}$ through a 2-to-1 multiplexer, allowing a reduction in the computational cost. However, applying this optimization directly to the design in Fig. 2(b) is less straightforward due to the exponential growth in the number of adders across stages in the parallel architecture. By strategically sharing partial sums across mirrored address locations (e.g., 0–3 and 4–7), the proposed shared LUT design effectively reduces the number of required computations by nearly 50%.
III-C3 Split LUT
In both parallel and shared LUT techniques, the number of adders and multiplexers grows exponentially. Although shared LUTs halve the computation, they still exhibit significant exponential scaling. To alleviate this, the LUT is split into two parts. The first $P$ terms in the summation defined by $\tilde{Y}_{r}$ in (8) can be implemented using the parallel LUT, requiring $(2^{P-1}-1)+(P-1)$ adders and a $2^{P-1}$-to-1 multiplexer (or equivalently, $2^{P-1}-1$ two-to-one multiplexers). The remaining terms are computed similarly, using $(2^{K-P-1}-1)+(K-P-1)$ adders and a $2^{K-P-1}$-to-1 multiplexer. One extra adder combines the two partial results. Thus, the total number of adders is $(2^{P-1}+2^{K-P-1}-2)+(K-2)$, and the total number of 2-to-1 multiplexers is $2^{P-1}+2^{K-P-1}-2$. A key observation is that decreasing $P$ reduces complexity in the first part but increases it in the second, and vice versa. By minimizing the total hardware with respect to $P$, it can be shown that the optimal configuration occurs at $P=\frac{K}{2}$. This implies the most efficient design is achieved when both partitions are of equal size. The proposed LUT architecture based on this technique is shown in Fig. 6(c). It can be observed that the chain of adders is replaced by a single adder for the first $K/2$ section, while the other term is computed using a subtractor. The remaining two partial sums are simply the two’s complements of these results. A similar structure applies to the second $K/2$ section as well.
III-C4 Hybrid LUT
Although the split LUT can reduce the computational challenges associated with parallel and shared LUTs, it still exhibits exponential dependence, which may be undesirable for larger values of $K$. To overcome this limitation, a hybrid LUT is proposed, as shown in Fig. 6(d), which pairs consecutive weights and utilizes minimal logic resources for the generation of partial sums. The term hybrid here refers to the transformation of large-scale parallelism into a smaller degree of parallelism, inspired by the serial-like pairing of weights. In this approach, adders and subtractors are used to compute the partial sums $\theta_{K-i+1}+\theta_{K-i}$ and $\theta_{K-i+1}-\theta_{K-i}$ for all $i\in\{1,2,\dots,K-2\}$, and can be shared as they are independent of selection lines. The last pair of weights can be generated using either a conditional adder or an adder/subtractor, depending on the design requirements. In this case, a conditional adder is employed to minimize resource usage. As a result, each pair of weights requires only one adder or subtractor, resulting in a linear scaling of resource requirements, which is significantly more efficient compared to other techniques. Furthermore, the selection lines for the 2-to-1 multiplexers—used to select between $\theta_{K-i+1}+\theta_{K-i}$ and $\theta_{K-i+1}-\theta_{K-i}$—can be derived from the observation that the contents at the $0^{\textnormal{th}}$ and $3^{\textnormal{rd}}$ address locations are two’s complements of each other, as are the $1^{\textnormal{st}}$ and $2^{\textnormal{nd}}$, and similarly for the second half, as illustrated in Fig. 2(a).

### v3

# COMET: Co-Optimization of CNN Models using Efficient-Hardware OBC Techniques

@@ Abstract
Convolutional Neural Networks (CNNs) achieve remarkable accuracy in vision tasks, yet their computational complexity challenges low-power edge deployment. In this work, we present COMET, a framework of CNN models that employ efficient hardware offset-binary coding (OBC) techniques to enable co-optimization of performance and resource utilization. The approach formulates CNN inference using OBC representations applied separately to inputs (Scheme A) and weights (Scheme B), enabling exploitation of bit-width asymmetry. The shift–accumulate operation is modified by incorporating offset-term with the pre-scaled bias. Leveraging symmetries in Schemes A and B, we introduce four look-up table (LUT) techniques—parallel, shared, split, and hybrid—and evaluate their efficiency. Building on this foundation, we develop a general matrix multiplication core using the im2col transformation for efficient CNN acceleration. We consider LeNet-5 and All-CNN-C to demonstrate that the OBC-GEMM core efficiently supports modern workloads. Evaluation shows that COMET enables efficient FPGA deployment compared to state-of-the-art designs, with negligible accuracy loss, demonstrating its efficiency and scalability across diverse network architectures.
@@
Convolutional Neural Network (CNN), Field-Programmable Gate Array (FPGA), General Matrix-Multiply (GEMM), Look-up Table (LUT), Offset-Binary Coding (OBC).
@@ I Introduction
Convolutional neural networks (CNNs) learn features directly from images and outperform traditional methods in classification, detection, and segmentation tasks [36]. Due to their strong representational capability, they are widely used in applications such as industrial defect inspection and medical imaging [18]. Modern CNNs incorporate deep topologies, residual and dense connections, attention mechanisms, normalization layers, depthwise separable convolutions, and multi-branch structures to improve performance [12]. However, this architectural complexity increases computational cost, as convolutions dominate the arithmetic workload in widely used architectures such as VGG-style [31], ResNet [28], MobileNet [26], and attention-augmented models [2].
Each convolution layer consists of numerous multiply–accumulate (MAC) operations, often reformulated using the im2col transformation into matrix–matrix or matrix–vector multiplications in the form inner product computations (IPCs) [6]. As network depth, channel count, and kernel size increase, these operations impose significant computational and memory demands, complicating deployment on low-power field programmable gate array (FPGA)-based platforms [37, 15]. Although FPGAs provide parallelism and reconfigurability, large MAC arrays quickly consume DSP resources, increase routing congestion, and raise power consumption, limiting scalability.
To improve resource utilization, many CNN accelerators adopt im2col-based general matrix multiplication (GEMM) architectures to reuse shared MAC arrays [6, 7]. While this approach improves throughput, it also increases DSP and memory usage. Small CNNs can be deployed relatively easily on modest hardware, whereas larger CNNs typically require bigger FPGA devices to accommodate the increased compute and storage demands. For resource-constrained devices, replacing DSPs with LUTs and BRAMs offers a promising direction, although the trade-off between LUT resource pressure and BRAM utilisation becomes increasingly important as network complexity grows. Even advanced dataflow strategies such as [19], which maximise data reuse, remain fundamentally dependent on MAC units, thereby constraining overall efficiency and scalability.
Distributed arithmetic (DA) is a multiplierless approach suited for low-power, area-constrained hardware [32, 13], replacing MAC operations in IPC with a pre-computed look-up table (LUT) and a shift-accumulate (SA) unit. This is achieved by representing the elements of a vector in an IPC either in two’s complement (TC) [16] or offset-binary coding (OBC) [24], with OBC halving the LUT size through symmetry. Recent DA-based LSTM optimizations use hardware LUTs [33, 17, 1, 14]. Parallel OBC-LUTs reduce critical path delay (CPD) but increase complexity [33], while serial LUTs with high-radix OBC/TC reduce memory but add latency [17, 14]. Few DA-specific CNN or 1-point convolution implementations exist [27, 4, 3], e.g., input pairing to shrink LUTs [27] or in-memory optimization to reduce memory [4]. For efficient CNN implementation on FPGA platforms, asymmetric quantization is effective: reducing the weight bit-width while keeping input precision fixed lowers power consumption at a modest accuracy cost [20]. However, this strategy has not yet been investigated in DA-based CNN architectures, particularly with respect to validation on FPGA platforms. Moreover, its impact on model accuracy and hardware efficiency across various CNN architectures has not been clearly established.
@@
To address these challenges, we propose COMET, our co-optimization framework for CNN hardware acceleration. For evaluation purposes, we adopt LeNet-5 [21] and All-CNN-C [29] as complementary benchmarks. LeNet-5, an early CNN with a compact structure and low parameter count, provides a controlled platform to analyze resource, performance, and accuracy trade-offs; it often requires lightweight modification for efficient deployment [9]. In contrast, All-CNN-C is a complex, modern, fully convolutional architecture, with strided convolutions replacing fully connected and pooling layers, stressing scalable IPCs without requiring modifications. Both networks rely on structured IPCs: dense layers use full weight matrices, and convolutional layers use sparse, weight-shared kernels. Through im2col, convolutions are reformulated as GEMM-based IPCs. COMET reduces DSP usage and power consumption without altering network functionality, employing a streaming (im2col + tiled) strategy and dynamic parallel LUTs to compute partial sums at runtime. Standard CNN kernels are executed as OBC-based GEMM operations within a scalable hardware architecture. Evaluation on LeNet-5 and All-CNN-C demonstrates that COMET generalizes across lightweight and convolution-dominated networks, validating both feasibility and scalability toward more complex CNNs. Our main contributions are as follows:
  • We propose a im2col-based formulation of the 2D CNN convolution layer using OBC representations of inputs (Scheme A) and weights (Scheme B), as initial part of the COMET framework.
  • We introduce four LUT architectures based on efficient hardware OBC techniques and optimize the SA unit by unifying pre-scaled bias and offset terms, improving hardware efficiency.
  • We explore and evaluate the design space of the OBC-GEMM framework using Scheme A and Scheme B based on the proposed LUT architectures, analyzing key design trade-offs.
  • We investigate the impact of fixed-point quantization on the accuracy of LeNet-5 and All-CNN-C models and integrate the proposed OBC-GEMM core into their convolutional kernels for hardware evaluation.
  • We compare the proposed architectures integrated with LeNet-5 and All-CNN-C against state-of-the-art hardware accelerators, highlighting the improvements in resource–energy trade-offs.
The remainder of this paper is organized as follows. Section II introduces the OBC formulation of a standard CNN. Section III presents the proposed LUT architectures and their co-optimization trade-offs. Section IV covers the evaluation of the OBC-GEMM module. Section V describes the CNN models consideration and accuracy results. Section VI provides comparison with state-of-the-art CNN accelerators, and Section VII concludes the paper.
Figure 2: OBC-based IPC for $K=4$ with: (a) Traditional LUT and SA-unit [24], (b) Hardware LUT (w/o pipelining) and CSFA-based SA unit [33].
@@ II CNN Formulation with OBC
As shown in Fig. 1, a standard CNN is a multi-layer pipeline where each layer transforms input feature maps of size $H\times W$ into higher-level representations using convolutional kernels of size $K\times L$ and biases. A layer with $C$ input and $N$ output channels computes each output by convolving all $C$ input channels with their respective kernels via a sliding-window operation (A) and summing across channels. This process is repeated for all $N$ outputs. For simplicity, a batch size of $1$ is assumed, though the formulation extends to multiple batches. Specifically, given an input tensor with $C$ channels, each with spatial dimensions $H\times W$, i.e., $X\in\mathbb{R}^{C\times H\times W}$, the output of a 2D convolution of $X$ with a filter tensor $\Theta\in\mathbb{R}^{N\times C\times K\times L}$ is a tensor $Y\in\mathbb{R}^{N\times(H+K)\times(W+L)}$, whose entries are
@@
where $n=1,\ldots,N$, $h=1,\ldots,H+K$, $w=1,\ldots,W+L$, and $\beta_{n}\in\mathbb{R}$ represents a bias term. Note  (1) describes the convolution computation for a single layer. Each product $X_{c,h+k,w+l}\Theta_{n,c,k,l}$ followed by its accumulation constitutes one MAC operation. Therefore, a single convolution layer requires roughly $N\times H\times W\times C\times K\times L$ MAC operations, and across multiple layers, total MACs can reach tens of millions, stressing compute and memory bandwidth [11]. To mitigate this, we adopt the im2col transformation [6], which converts convolutions into tile-based GEMM, enabling the reuse of a single MAC array across layers. Applied to (1), the transformation typically follows these steps:
@@
  • The convolution is expressed as an IPC of length $N_{p}=CKL$ between the flattened input patch and filter:
@@
The resulting output is then reshaped into the proper order. For simplicity, we set $C=1$ and $L=1$ in the IPC formulation, though the analysis extends to any $C$ and $L$. Denoting the summation in (2) as $\tilde{Y}$}, we have:
@@
where $x_{i,r}\in\{0,1\}$ is the $r$th bit of $x_{i}$. In OBC representation, $x_{i}$ can be written as $x_{i}=\frac{1}{2}[x_{i}-(-x_{i})]$, which results (4) in
@@
where $\bar{x}_{i,r}$ is the ones complement of $x_{i,r}$ and $\lfloor\cdot\rfloor$ is the floor function. Substituting (5) and (6) into (3), we have
@@
Each LUT entry represents a combination of weights with input bit-slices $(x_{n,B_{1}-1},\ldots,x_{n,0};1\leq n\leq K)$, ordered from least to most significant bit, with bit-slices used as addresses at runtime. Outputs are accumulated via a SA operation over $B_{1}$ clock cycles to compute $\tilde{Y}$, as shown in Fig. 2(a). The offset term $\theta_{ot}$ is the twos complement of the LUT content at address 0, loaded at the start of shift-accumulation, as shown in Fig. 2(a). Traditional LUTs become inefficient for large $K$ due to access time. Dynamic generation using adders and multiplexers [33] lacks scalability and adds hardware overhead: $2^{K}(1-2^{-K/2})$ adders/subtractors, $2^{K-1}-1$ 2-to-1 multiplexers, and CPD of $\log_{2}K\cdot T_{A}+(K-1)\cdot T_{MX}$, where $T_{A}$ and $T_{MX}$ are adder and multiplexer delays. Note carry-save full adders (CSFAs) based SA unit has been employed to reduce this delay.
Following (11), $Y$ can be computed using Scheme B, where each weight $\theta_{i}$ in OBC form with bit-width $B_{2}$ is processed as in (4)–(11), yielding:
@@
Note that Scheme B shares the same structure as Scheme A, except that the LUT content and SA clock cycles may differ ($B_{1}\neq B_{2}$). The schemes differ only in which component is represented in OBC form, leveraging the weight–input bit-width asymmetry to offer hardware implementation trade-offs.
@@ III-A System Overview
At the start of each layer, a configuration word is loaded, encoding kernel dimensions ($C\times K\times L$), stride ($S$, with $S=1$ for normal and $S=2$ for downsampling), padding type/amount ($P$, with $P=0$ for no padding and $P=1$ for one-sided zero padding), input/output channels (CC/NN), and bit-width ($B_{1}$ or $B_{2}$). A cycle counter drives the im2col address generator, computing read addresses dynamically and injecting zeros for padding. Inputs are fetched from RAM and streamed into the OBC-GEMM engine, while the next chunk is prefetched via ping–pong buffers. Partial or final sums are written back to RAM at the generated addresses, ensuring correct spatial alignment. This repeats for all layers, forming a fully streamed, stall-free pipeline.
@@ III-B Address Generator and Control Unit
At the core of the address generator is a set of nested counters managed by the FSM, handling address generation and read/write scheduling. As shown in Fig. 4, the architecture includes a dedicated cntr0 and three hierarchical counter groups for read (rd_cntr#$m$), calculation (cal_cntr#$m$), and write (wr_cntr#$m$) operations. Each group has four levels of counters indexed by $m\in\{1,4\}$, corresponding to tiling, spatial position $(h,w)$, output channel $(n)$, and layer index. cntr0 tracks the OBC-GEMM clock cycle. Signals 14 in Fig. 4 indicate the hierarchical carry mechanism, enabling stepwise progression and value transfer between read, calculation, and write stages.
In this hierarchical carry mechanism, rippling primarily affects the read counters. The clock cycle counter (cntr0) generates sequential memory addresses for one tile-level OBC-GEMM operation. When it reaches its upper bound, all data for the current tile are loaded, carry 1 is asserted, cntr0 resets, and rd_cntr1 (tile index) increments to start the next tile. When rd_cntr1 completes, carry 2 resets it and increments rd_cntr2 (spatial position). Similarly, 3 resets rd_cntr2 and increments rd_cntr3 (output channel), and 4 resets rd_cntr3 and increments rd_cntr4 (layer index) to begin the next layer.
@@
III-C OBC-GEMM Core
The OBC-GEMM core comprises $L$ LUTs of size $K$, $L$ SA units, and $L$ PISO units of size $K$ for handling input or weight bit-slices, as shown in Fig. 5. LUTs dynamically generate partial sums via adders and multiplexers, while SA units combine them efficiently. PISO units serialize inputs or weights to match the dataflow of Scheme A or B. A key optimization is a single adder for both offset and bias computation, eliminating separate bias logic and reducing hardware complexity, area, and power.
III-C1 Parallel LUT
This initial architecture uses fewer adders than the one shown in Fig. 2(b) and generates all LUT contents from the value at address 0. Chain adders reuse the partial sum $\theta_{3}+\theta_{4}$, which appears in the LUTs at addresses 3, 4, and 7, as illustrated in Fig. 6(a). This sum is added to $\theta_{1}$, present in all addresses, and then added or subtracted from $\theta_{2}$ to produce all OBC combinations. The LUT at address 1 can be obtained by subtracting $2\theta_{2}$ (via a right shift) from address 0. This process continues with subtractors for all addresses. While CPD grows linearly with $K$, the number of adders and multiplexers grows exponentially. To limit this, $K$ can be factored as $K=p\times q$, restricting the exponential growth to $q$, with higher-order $K$ implemented using an adder tree of depth $\log_{2}p$.